OpenStudy (anonymous):

please help! Write the equation of a hyperbola with vertices (0,-2) and (0,2) and co-vertices (-4,0) and (4,0)

OpenStudy (anonymous):

Answers are: \[\frac{ y ^{2} }{ 16 }-\frac{ x ^{2} }{ 4 }=1\] \[\frac{ x ^{2} }{ 4 }-\frac{ y ^{2} }{ 16 }=1\] \[\frac{ y ^{2} }{ 4 }-\frac{ x ^{2} }{ 16 }=1\]

OpenStudy (anonymous):

answer is last one. because x=-4 and y=0 is true in that option

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

can you help me ith another one?

OpenStudy (anonymous):

yes of course :)

OpenStudy (anonymous):

Write the equation of a hyperbola with vertives (0,-4) and (0,4) and foci (0,-5) and (0,5)

OpenStudy (anonymous):

\[\frac{ y ^{2} }{ 9 }-\frac{ x ^{2} }{ 16 }=1\] \[\frac{ x ^{2} }{ 9 }-\frac{ y ^{2} }{ 16 }=1\] \[\frac{ y ^{2} }{ 16 }-\frac{ x ^{2} }{ 9 }\]

OpenStudy (anonymous):

does last one equal 1? if it does, answer is that

OpenStudy (anonymous):

oh yeah lol sorry.

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

you're welcome :)